OpenStudy (anonymous):

use logarithm differentiation to find dy/dx for the following y=(8x + 1)^3 e^x/ sqrt(4x^2 + 1)

OpenStudy (anonymous):

take ln of both sides \[lny = \ln[(8x+1)^3]+lne^x-\ln[\sqrt{4x^2+1}]\]

OpenStudy (anonymous):

\[lny = 3\ln[8x+1]+xlne-1/2\ln[4x^2+1]\]

OpenStudy (anonymous):

\[lny = 3\ln[8x+1]+x-1/2\ln[4x^2+1]\]

OpenStudy (anonymous):

lne = 1 so the lne goes away in that last one

OpenStudy (anonymous):

\[1/y*dy/dx = 24/(8x+1)+1-4/(4x^2+1)\]

OpenStudy (anonymous):

\[1/y*dy/dx = 24/[4*(2x+1)]+1-4/[(2x+1)(2x-1)]\]

OpenStudy (anonymous):

etc multiply by y and replace y with the original equation